In an experiment to measure Young’s modulus for steel, a wire is suspended vertically and loaded at the free end. In such an experiment, (a) why is the wire long and thin (b) why is a second steel wire suspende adjacent to the first? Sketch the graph you will expect to obtain in such an experiment showing the relation between the applied load and the extension of the wire. Show how it is possible to use the graph to determine (a) Young’s modulus for the wire(b) the work done in stretching the wire.
1
Expert's answer
2020-11-02T09:26:09-0500

A) A wire long and thin that stretching was more noticeable.

B) To apply force for stretching of the first wire.

 


A)As Jung's module is equal to "E=\frac {F\times\Delta L }{S\times L}" therefore it is conveniently possible to find numerator.

B) As work is equal to "A=F\times \Delta L" , too it is convenient to find it according to the schedule.